A basket contains 80 mangoes and 60 oranges ,if two fruit are picked one after the other without replacement ,what is the probability that (i) one of each fruit is picked (ii)one type of fruit is picked?
Answer by Edwin McCravy(20056)

A basket contains 80 mangoes and 60 oranges, if two fruit
are picked one after the other without replacement, what is
the probability that
(i) one of each fruit is picked
P[(mango then orange) OR (orange then mango)]

(80/140)*(60/139) + (60/140)*(80/139] = 480/973

(ii)one type of fruit is picked?
Method 1 [using (i), easier]: 

subtract 1 - 480/973 = 493/973

Method 2 [from scratch, harder]: 

P[(mango then mango) OR (orange then orange)]

(80/140)*(79/139) + (60/140)*(59/139] = 493/973 

Edwin